OpenStudy (anonymous):

Find the x- and y-intercepts and then need to graph y=3/2x-3

OpenStudy (anonymous):

intercepts are when the other coordinate =0 (x,0) is the x-intercept and (0,y) is the y-intercept y=mx+b, b is the y-intercept so (0,-3) is the y-intercept when y=0. 0=3/2x-3, 3=3/2x, 6=3x, x=2, (2,0) is the x-intercept

OpenStudy (anonymous):

thank you